OpenStudy (anonymous):

Which equation represents the line perpendicular to the line x + 2y = 12?

OpenStudy (anonymous):

A. y = -2x -9 B. y = (1/2)x + 6 C. 2x - y = -3 D. 2x + 4y =7

OpenStudy (anonymous):

slope of line y=m x+c is m and slope of line perpendicular to it is -1/m

OpenStudy (domebotnos):

First solve to find the equation in point-slope form. x + 2y = 12 -x -x 2y = -x + 12 __ ___ ____ 2 2 2 y = -1/2 + 6

OpenStudy (domebotnos):

Then flip it to find the slope of the perpendicular line. -1/2 flipped and the sign reversed is 2. That narrows it down to C. or D.

OpenStudy (anonymous):

so its B

OpenStudy (domebotnos):

Look at how you solve D. 2x + 4y = 7 -2x -2x 4y = -2x + 7 __ ____ ___ 4 4 4 y = -1/2 + 7/4 Now look at C. 2x - y = -3 -2x -2x -y = -2x - 3 ___ ____ ___ -1 -1 -1 y = 2x + 3 The answer is C.

OpenStudy (anonymous):

Thank you both very much